OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–3, 7); Slope: 4 A. y – 7 = 4(x – 3) B. y – 7 = 4(x + 3) C. y + 7 = 4(x + 3) D. y + 7 = 4x – 12

OpenStudy (texaschic101):

point slope form is : y - y1 = m(x - x1) slope(m) = 4 (-3,7) x1 = -3 and y1 = 7 now just sub y - y1 = m(x - x1) y - 7 = 4(x - (-3) y - 7 = 4(x + 3) <=== here you go, in point slope form

OpenStudy (anonymous):

thnx @texaschic101

OpenStudy (texaschic101):

no problem :)